# Beyond Maximum Likelihood: Variational Inequality Estimation for Generalized Linear Models

### v1

@@ Appendix C Proofs for Section 4
$\mu\|\bm{\beta}^{\star}-\hat{\bm{\beta}}_{N}\|\leq\|V_{N}(\bm{\beta}^{\star})\|.$ (C.1)
@@
Then we know from Lemma 4.2 that
@@
As $V(\bm{\beta}^{\star})=\mathbf{0}$, by applying the Hoeffding’s inequality [vershynin2018high, Theorem 2.2.6], it follows that for any $t>0$ and $j\in\{0,\dots,d\}$,
@@
Combining this with (C.1), we derive the desired results. ∎
@@
$\frac{V(\bm{\beta}^{\star}+t\bm{h}_{t})-V(\bm{\beta}^{\star})}{t}=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\left[\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\right$]. (C.2)
@@
$\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\ \rightarrow\ (g^{-1})^{\prime}(u)\,\langle\tilde{\bm{x}},\bm{h}\rangle\,\tilde{\bm{x}}$. (C.3)
@@
Under Assumption 4.1, we have $\|\tilde{\bm{x}}\|^{2}\leq 1+dM^{2}$, so the right-hand side is bounded by $L(1+dM^{2})\|\bm{h}_{t}\|$, which is integrable and independent of $t$ since $\|\bm{h}_{t}\|$ is bounded. From (C.2) and (C.3) with the dominated convergence theorem, it follows that
@@
$\sqrt{N}\big(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star})\big)\ \xrightarrow{d}\ \mathcal{N}(\mathbf{0},\,\Gamma).$ (C.4)
@@
$\mathbf{0}=V_{N}(\hat{\bm{\beta}}_{N})$ $=V_{N}(\bm{\beta}^{\star})+\underbrace{V(\hat{\bm{\beta}}_{N})-V(\bm{\beta}^{\star})}_{\text{population increment}}+\underbrace{(V_{N}(\hat{\bm{\beta}}_{N})-V(\hat{\bm{\beta}}_{N}))-(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star}))}_{=:R_{N}}.$ (C.5)
@@
$V(\hat{\bm{\beta}}_{N})-V(\bm{\beta}^{\star})=\nabla V(\bm{\beta}^{\star})(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})+r_{N},\qquad\|r_{N}\|=o(\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|).$ (C.6)
From Lemma 4.2, the sample operator $V_{(\bm{x},y)}$ is $(LdM^{2}+L)$-Lipschitz. Therefore,
@@
Each function $f_{\bm{\beta}}$ is uniformly bounded by the envelope $LdM^{2}+L$, and the class $\mathcal{F}_{r}$ is a finite-dimensional (Euclidean) Lipschitz image of a ball in $\mathbb{R}^{d+1}$. Such finite-dimensional Lipschitz (parametric) classes have a covering numbers that grow polynomially in $1/\epsilon$, a direct consequence of the parameter space being a subset of $\mathbb{R}^{d+1}$ (see [van1996weak, Theorem 2.7.11]). This polynomial entropy bound, together with the existence of a square-integrable envelope, implies that $\mathcal{F}_{r}$ is $\mathbb{P}$-Donsker by [van1996weak, Theorem 2.5.2]. Consequently,
$\sup_{f\in\mathcal{F}_{r}}\big\|\sqrt{N}[\mathbb{E}_{N}-\mathbb{E}]f\big\|=O_{p}(1).$ (C.7)
@@
This together with $\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|=O_{p}(N^{-1/2})$ and (C.7), we conclude on $\mathcal{A}_{N}$ that
@@
The first term converges in distribution to $\mathcal{N}(0,\nabla V(\bm{\beta}^{\star})^{-1}\Gamma\nabla V(\bm{\beta}^{\star})^{-\top})$ by (C.4). On the other hand, $\sqrt{N}\|r_{N}\|=o_{p}(1)$ and $\sqrt{N}\|R_{N}\|=o_{p}(1)$, the second term vanishes in probability. Applying Slutsky’s theorem gives
@@
Appendix D Proofs for Section 5
@@
By the strong Minty monotonicity condition in Assumption 4.1 and the Lipschitz continuity in Lemma 4.2, we have
@@
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t+1}-\bm{\beta}^{\star}\|^{2}$ $\leq\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}-\eta^{t}\,\mathbb{E}\big[\langle V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t}),\bm{\beta}^{t}-\bm{\beta}^{\star}\rangle\big]+\frac{1}{2}(\eta^{t})^{2}\,\mathbb{E}\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}.$ (D.1)
@@
In addition, Lemma 4.2 implies that $\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}\leq(d+1)R^{2}M^{2}$. Substituting these bounds into (D.1) yields
@@
Appendix E Additional Experiment Results

### v2

@@ Appendix D Proofs for Section 4
$\mu\|\bm{\beta}^{\star}-\hat{\bm{\beta}}_{N}\|\leq\|V_{N}(\bm{\beta}^{\star})\|.$ (D.1)
@@
Then we know from Lemma D.1 that
@@
As $V(\bm{\beta}^{\star})=\mathbf{0}$, by applying the Hoeffding’s inequality [22, Theorem 2.2.6], it follows that for any $t>0$ and $j\in\{0,\dots,d\}$,
@@
Combining this with (D.1), we derive the desired results. ∎
@@
$\frac{V(\bm{\beta}^{\star}+t\bm{h}_{t})-V(\bm{\beta}^{\star})}{t}=\mathbb{E}_{(\bm{x},y)\sim\mathbb{P}}\left[\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\right$]. (D.2)
@@
$\frac{g^{-1}(u+t\delta_{t})-g^{-1}(u)}{t}\,\tilde{\bm{x}}\ \rightarrow\ (g^{-1})^{\prime}(u)\,\langle\tilde{\bm{x}},\bm{h}\rangle\,\tilde{\bm{x}}$. (D.3)
@@
Under Assumption 4.1, we have $\|\tilde{\bm{x}}\|^{2}\leq 1+dM^{2}$, so the right-hand side is bounded by $L(1+dM^{2})\|\bm{h}_{t}\|$, which is integrable and independent of $t$ since $\|\bm{h}_{t}\|$ is bounded. From (D.2) and (D.3) with the dominated convergence theorem, it follows that
@@
$\sqrt{N}\big(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star})\big)\ \xrightarrow{d}\ \mathcal{N}(\mathbf{0},\,\Gamma).$ (D.4)
@@
$\mathbf{0}=V_{N}(\hat{\bm{\beta}}_{N})$ $=V_{N}(\bm{\beta}^{\star})+\underbrace{V(\hat{\bm{\beta}}_{N})-V(\bm{\beta}^{\star})}_{\text{population increment}}+\underbrace{(V_{N}(\hat{\bm{\beta}}_{N})-V(\hat{\bm{\beta}}_{N}))-(V_{N}(\bm{\beta}^{\star})-V(\bm{\beta}^{\star}))}_{=:R_{N}}.$ (D.5)
@@
$V(\hat{\bm{\beta}}_{N})-V(\bm{\beta}^{\star})=\nabla V(\bm{\beta}^{\star})(\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star})+r_{N},\qquad\|r_{N}\|=o(\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|).$ (D.6)
From Lemma D.1, the sample operator $V_{(\bm{x},y)}$ is $(LdM^{2}+L)$-Lipschitz. Therefore,
@@
Each function $f_{\bm{\beta}}$ is uniformly bounded by the envelope $LdM^{2}+L$, and the class $\mathcal{F}_{r}$ is a finite-dimensional (Euclidean) Lipschitz image of a ball in $\mathbb{R}^{d+1}$. Such finite-dimensional Lipschitz (parametric) classes have a covering numbers that grow polynomially in $1/\epsilon$, a direct consequence of the parameter space being a subset of $\mathbb{R}^{d+1}$ (see [21, Theorem 2.7.11]). This polynomial entropy bound, together with the existence of a square-integrable envelope, implies that $\mathcal{F}_{r}$ is $\mathbb{P}$-Donsker by [21, Theorem 2.5.2]. Consequently,
$\sup_{f\in\mathcal{F}_{r}}\big\|\sqrt{N}[\mathbb{E}_{N}-\mathbb{E}]f\big\|=O_{p}(1).$ (D.7)
@@
This together with $\|\hat{\bm{\beta}}_{N}-\bm{\beta}^{\star}\|=O_{p}(N^{-1/2})$ and (D.7), we conclude on $\mathcal{A}_{N}$ that
@@
The first term converges in distribution to $\mathcal{N}(0,\nabla V(\bm{\beta}^{\star})^{-1}\Gamma\nabla V(\bm{\beta}^{\star})^{-\top})$ by (D.4). On the other hand, $\sqrt{N}\|r_{N}\|=o_{p}(1)$ and $\sqrt{N}\|R_{N}\|=o_{p}(1)$, the second term vanishes in probability. Applying Slutsky’s theorem gives
@@
Appendix E Comparison of Asymptotic Efficiency between VI and MLE
shares the same “sandwich” structure as that of the MLE,
where the expected Hessian and the score covariance are given by
and
[…]
For correctly specified models, the information identity
$\Sigma_{\mathrm{MLE}}\ \preceq\ \Sigma_{\mathrm{VI}}.$
Appendix F Proofs for Section 5
@@
By the strong Minty monotonicity condition in Assumption 4.1 and the Lipschitz continuity in Lemma D.1, we have
@@
$\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t+1}-\bm{\beta}^{\star}\|^{2}$ $\leq\frac{1}{2}\mathbb{E}\|\bm{\beta}^{t}-\bm{\beta}^{\star}\|^{2}-\eta^{t}\,\mathbb{E}\big[\langle V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t}),\bm{\beta}^{t}-\bm{\beta}^{\star}\rangle\big]+\frac{1}{2}(\eta^{t})^{2}\,\mathbb{E}\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}.$ (F.1)
@@
In addition, Lemma D.1 implies that $\|V_{(\bm{x}^{t},y^{t})}(\bm{\beta}^{t})\|^{2}\leq(d+1)R^{2}M^{2}$. Substituting these bounds into (F.1) yields
@@
Appendix G Additional Experiment Results
G.1 GLM experiment
@@ VI MLE VI MLE VI MLE VI MLE
G.2 GAM experiment
((a)) log
((b)) softplus
((c)) clipped exp.
((d)) GMM CDF
Figure 4: Reconstruction of additive components in GAMs under Poisson distribution.
((a)) logistic
((b)) arctangent
Figure 5: Reconstruction of additive components in GAMs under Bernoulli distribution.
((a)) log
((b)) softplus
((c)) clipped exp.
((d)) GMM CDF
Figure 6: Squared error of parameter estimates and the average $L^{2}$ error of the additive components (Poisson response)
((a)) logistic
((b)) arctangent
Figure 7: Squared error of parameter estimates and the average $L^{2}$ error of the additive components (Bernoulli response)
We report detailed experimental settings and results for GAMs considered in Section 7. Each additive component $f_{j}$ is represented using a truncated Legendre basis of degree $K$ without the constant term, so that $f_{j}(0)=0$ holds by construction. We use $d=4$ additive components and $K=3$ basis functions per component, with covariates independently sampled from U[1,1]U[-1,1]. Each $\beta_{jk}$ is drawn from a standard normal distribution and then normalized so that $\|\bm{\beta}\|_{2}=1$. We consider the Poisson and Bernoulli response models.
For Poisson models, we consider the canonical log link as well as softplus, clipped exponential, and Gaussian-mixture CDF links, as in the GLM experiments of Section 6. For Bernoulli models, we consider the canonical logistic link together with the non-canonical arctangent link, whose inverse is given by
$g^{-1}(z)=\frac{1}{2}+\frac{1}{\pi}\arctan(z).$
Figures 4 and 5 compare the recovered additive components with the ground truths, while Figures 6 and 7 report the squared error of the parameter estimates and the average $L^{2}$ errors of the additive components over iterations for the Poisson and Bernoulli models, respectively. For non-canonical links in both Poisson and Bernoulli GAMs, the VI estimator often yields slightly smoother optimization trajectories and modestly improved reconstruction of the additive components compared to MLE, although the final errors are typically of similar magnitude. These experiments illustrate that the proposed VI framework extends naturally to GAMs through basis representations and remains numerically stable in additive models beyond linear predictors.